OpenStudy (anonymous):

What is the solution to the equation 3 1 + 2x = 2 ? What is the solution to the equation 3 1 + 2x = 2 ? @Mathematics

OpenStudy (anonymous):

\[3^1+2x=2\]

OpenStudy (anonymous):

^(1+2x)

OpenStudy (anonymous):

wait whats the equation?

OpenStudy (anonymous):

3^(1+2x)=2

OpenStudy (anonymous):

3^(1+2x) = 2 I reckon

OpenStudy (anonymous):

Alright so heres how you do it...

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

3^(1+2x) = 2 Take the log of both sides: log[3^(1+2x)]=(1+2x)log(3) (thats just a property of log) So now we have: (1+2x)log(3)=log(2) Distribute (1+2x) out: log(3)+2x(log(3))=log(2) And then just solve for x: 2x(log(3))=log(2)-log(3) x=(log(2)-log(3))/(2log(3)) <----- Thats the answer

OpenStudy (anonymous):

my options are x=0.292.....x=-0.185...x=0.185...x=0.292

OpenStudy (anonymous):

jp's answer is right. To get digits you'll need to use a calc.

OpenStudy (anonymous):

Alright... then its -0.185

OpenStudy (anonymous):

I though you needed the answer in exact form